OpenStudy (anonymous):

how do you evaluate like like variables with exponents? such as, 3^-4 * 3^4

terenzreignz (terenzreignz):

No difference. Remember \[\Large a^ma^n = a^{m+n}\] It always applies, even with negative exponents.

OpenStudy (anonymous):

oh ok, and if i were to simplify the product i would just find the reciprocal, right?

terenzreignz (terenzreignz):

First, what's your product in its 'raw' form (ie, just add their exponents)

OpenStudy (anonymous):

i'm trying to figure it out. the problem is to evaluate 3^-3 * 3^4 * 3 * 3^-5

OpenStudy (anonymous):

i know how to add the exponents but how do i incorporate that into multiplying the bases

terenzreignz (terenzreignz):

You don't multiply bases :) If they're the same base, you just copy the base. Let me illustrate: \[\Large \color{red}{a^2}\cdot\color{blue}{ a^4}\\\Large =\color{red}{a\cdot a}\cdot \color{blue}{a\cdot a\cdot a\cdot a}\\\Large = a^6\\\Large = a^{\color{red}2+\color{blue}4}\]

OpenStudy (anonymous):

i know that but these bases are numbers. does that principle still count the same for variables and numbers?

terenzreignz (terenzreignz):

Yup :) Variables are just placeholders for numbers :D

OpenStudy (anonymous):

the answers dont have a 3 in them D:

OpenStudy (anonymous):

omg algebra >.<

terenzreignz (terenzreignz):

Okay, first tell me the answer in its raw form. Just the 3 and its exponent... what is it?

OpenStudy (anonymous):

3^-3 ?

terenzreignz (terenzreignz):

<ahem> -3 , 4 , -5 ???

terenzreignz (terenzreignz):

Could you add these up again? :3

OpenStudy (anonymous):

is the answer 

OpenStudy (anonymous):

;-;

terenzreignz (terenzreignz):

I don't know. Check it. What is the answer in its raw form?

OpenStudy (anonymous):

-81 ?..

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

ugh

terenzreignz (terenzreignz):

-81? I don't even...

OpenStudy (anonymous):

i feel really stupid right about now

terenzreignz (terenzreignz):

Come on, you were close. I told you: Copy the base, and its exponent would be the sum of all the exponents in the factors.

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

YES. right?..

terenzreignz (terenzreignz):

No.

OpenStudy (anonymous):

shame

terenzreignz (terenzreignz):

\[\Large 3^{-3}\cdot 3^4\cdot 3^{-5}\]

terenzreignz (terenzreignz):

Step by step... copy the base... what's the base?

OpenStudy (anonymous):

waitwaitwaitwaitwait

terenzreignz (terenzreignz):

waiting

OpenStudy (anonymous):

3^-3 * 3^4 * 3 * 3^-5

OpenStudy (anonymous):

that was the problem

terenzreignz (terenzreignz):

Oh. Well, in that case, all correct :P Except...

terenzreignz (terenzreignz):

Sorry for that mix-up :3

OpenStudy (anonymous):

what about the lone three? since it has no exponent doesn't that automatically give it an exponent of 1?

terenzreignz (terenzreignz):

Yes, it does.

OpenStudy (anonymous):

so i was right ._.

terenzreignz (terenzreignz):

\[\Large 3^{-3}\] Then, and you're correct up to this point.

OpenStudy (anonymous):

oh my god

OpenStudy (anonymous):

im so done

terenzreignz (terenzreignz):

Something wrong?

terenzreignz (terenzreignz):

Form the looks of it, that's not your final answer yet.

OpenStudy (anonymous):

XD thanks for youe help. i know more on how to work these kinds of problems out

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

you're kidding

terenzreignz (terenzreignz):

Well, is \(\Large 3^{-3}\) one of your choices?

OpenStudy (anonymous):

no i was thinking it was 1/81

terenzreignz (terenzreignz):

It isn't.

OpenStudy (anonymous):

okay well my other answer is 1/27

terenzreignz (terenzreignz):

You can't have two different answers. I want you to be sure. Next rule on exponents. \[\Large a^{-n}\]

terenzreignz (terenzreignz):

When you have a negative exponent, it may be brought down to the denominator AND THEN replaced with a positive exponent. \[\Large a^{-n} = \frac1{a^n}\]

terenzreignz (terenzreignz):

What does that bode for \[\Large 3^{-3}=\color{red}?\]

OpenStudy (anonymous):

yes. it was -27 then i turned it to 1/27. okay this answer is right. i know it is

terenzreignz (terenzreignz):

It's not -27 :/ -27 is not the same as 1/27. Better write it as \(\Large 27^{-1}=\frac1{27}\)

OpenStudy (anonymous):

k

terenzreignz (terenzreignz):

So much for that.

OpenStudy (anonymous):

he answers are a.-27 b.1/27 c.-81 d.1/81

OpenStudy (anonymous):

it's b

terenzreignz (terenzreignz):

Show me that you understand how to do this... What would be the fraction form of \[\Large 5^{-2}=\color{red}?\]

OpenStudy (anonymous):

-10?...

terenzreignz (terenzreignz):

Nope. I told you... \[\Large a^{-n}= \frac1{a^n}\]

OpenStudy (anonymous):

ugh i feel like a complete idiot.

OpenStudy (anonymous):

i'm trying to understand all of this but it's difficult

terenzreignz (terenzreignz):

Okay. Let's start at square one again... \[\Large a^{-n}=\]

OpenStudy (anonymous):

1/a^n

terenzreignz (terenzreignz):

Okay... in that case, what's \[\Large 5^{-2}=\color{red}?\]

OpenStudy (anonymous):

i dont know...

terenzreignz (terenzreignz):

That's why you should not jump straight to the answer and ignore the process -.- Now... \[\Large a^{-n }=\]

terenzreignz (terenzreignz):

When you have a negative exponent, you bring the base into the denominator... \[\Large a^{-n}= \color{red}{\frac1{a^?}}\]

OpenStudy (anonymous):

look dude i really appreciate your help but i clearly dont have the mental capacity to understand this stuff. i'm just going to go ask a physical person for help. that usually does better for me. thanks anyways mate.

terenzreignz (terenzreignz):

And its new exponent would be its previous exponent, multiplied by -1. \[\Large a^{-n}=\frac1{a^{-n\cdot-1}}=\frac1{a^n}\]